how far from the base pf a 24 unit wall should a 26-unit ladder be leaning so that it reaches the top most part of the wall


1
Expert's answer
2022-03-29T17:51:37-0400

By Pythagorean Theorem "a^2+b^2=c^2"

Here "a=24,b=??, c=26"

"24^2+b^2=26^2"

"b^2=26^2-24^2" =100

"b=\sqrt{100}=10" unit